OpenStudy (anonymous):

A nut mixture is made up of 24 lbs of less expensive nuts that cost $1.80 per pound and a more expensive kind of nut (at $3.80 per pound). If the cost of the resulting mixture is $ 2.60 per pound then how many pounds of the expensive nut were added?

OpenStudy (amistre64):

x lbs + y lbs = total lbs x lbs($) + y lbs($) = (x+y lbs)(cost)

OpenStudy (amistre64):

24(1.80) + exp(3.80) = (24 + exp)(2.60)